OpenStudy (anonymous):

Which is the simplified form of 6 to 2? 1 to 4 one-third 3 to 1 four colon one

OpenStudy (calculusxy):

Find the greatest common factor.

OpenStudy (calculusxy):

So you know the GCF is 2 itself. So you would have to divide 6/2=3 and 2/2=1. So the reduced term will be 3/1.

OpenStudy (anonymous):

oh ok ok thank you so much :) i still have a couple of other questions i'm having trouble with though i'll ask someone else!

OpenStudy (anonymous):

3 to 1

OpenStudy (mathstudent55):

\(\dfrac{6}{2} \) \(= \dfrac{2 \times 3}{2 \times 1} \) \(= \dfrac{2}{2} \times \dfrac{3}{1} \) \(= 1 \times \dfrac{3}{1} \) \(= \dfrac{3}{1} \)